OpenStudy (anonymous):

please help me factor: (D^3 - 3D^2 + 4)

OpenStudy (lgbasallote):

*sigh* factoring differential equations =_= i always hated these stuffs :/

OpenStudy (anonymous):

yuuuup!

OpenStudy (anonymous):

how would (D-2)^2(D+1)?

OpenStudy (anonymous):

sound*?

OpenStudy (lgbasallote):

that would be (D-2)(D-2) = D^2 - 2D + 4 (D^2-2D+4)(D+1) = D^3 -2D^2 + 4D + 1 right?

OpenStudy (anonymous):

-4D in top part, right?

OpenStudy (lgbasallote):

D(2) = 0 right? unless you're doing this algebraically..

OpenStudy (anonymous):

even so, isnt (d-2)^2 = d^2 - 4d +4?